OpenStudy (kainui):

This is bugging me, I feel like there's gotta be some simple answer but I can't seem to get it.

OpenStudy (kainui):

Show that if \(n\) is odd, then: \(\gcd(n,n+2)=1\)

OpenStudy (dan815):

oh i got it

OpenStudy (dan815):

umm okay how about we just think about in multiple sense

OpenStudy (dan815):

if its the the multiple of some prime u know that there has to be that prime difference atleast

OpenStudy (dan815):

nd the only prime difference with 2 differences are 2, and this is odd so

OpenStudy (dan815):

does that make sense?

OpenStudy (dan815):

so for example if u have 15 and 17 now 3 ,5 divide 15 but u know 3 cannot divide 17 because there a difference of 2 and 5 cannot dice it because again a diff of 2

OpenStudy (dan815):

similarly all the primes that divide some odd number cannot divide another odd number just 2 away, as they shud all have remainders building up as all the odd primes 3 or more

OpenStudy (dan815):

just think about multiples of primes

OpenStudy (dan815):

theres no way another multiple of the same prime will exist just 2 away

OpenStudy (dan815):

for odd ofc

OpenStudy (dan815):

GCD(n,n+2)=2, if n is even?

OpenStudy (dan815):

yup

OpenStudy (kainui):

I think I get what you're saying, like since the gap is smaller than 3 (the first prime they could share in common) then they have to be relatively prime? I guess this makes sense but I don't know if I'm 100% convinced for some reason haha.

OpenStudy (freckles):

\[\gcd(n,n+2)=\gcd(n,2)=1,2, |n| \\ \text{ but } 2 \text{ is not a factor of } n \\ \text{ so } \gcd(n,2) \neq 2 \\ \\ \text{ in order for } \gcd(n,2)=|n| \\ \text{ we would have to have } \\ \text{ that } |n| \text{ is a factor of } 2 \\ \text{ which means } |n|=1 \text{ or } |n|=2 \\ \text{ but we already said } n \neq 2 \\ \text{ so } |n|=1 \]

OpenStudy (dan815):

yah

OpenStudy (dan815):

like it has to be true

OpenStudy (dan815):

think about it

OpenStudy (kainui):

Ok here I got a way of showing it that convinces myself: Suppose \(p\) divides \(n\) and \(n+2\) then that means I can write \(n=pa\) and \(n+2=pb\) so: \[n+2=n+2\] \[pa+2=pb\] that means \(p=2\) is the only prime that works, cause 2 has to also be divisible by p.

OpenStudy (dan815):

all other prime multiple of the other number will generate a remainder

OpenStudy (dan815):

okay yep thats pretty much same thing

OpenStudy (dan815):

so now u know there has to be this remainder

OpenStudy (dan815):

as p is greather than 2

OpenStudy (kainui):

Oh wow ok I should be thinking like @freckles is, I didn't even think about this: \[\gcd(n,n+2)=\gcd(n,2)\] ok in terms of remainders and the Euclidean algorithm this makes sense

OpenStudy (dan815):

oh i can see why that formula is true hehe :D

OpenStudy (kainui):

thanks everyone :D

OpenStudy (dan815):

pls medul

OpenStudy (dan815):

+fAN

OpenStudy (freckles):

hmmm... to medal dan or to not

OpenStudy (freckles):

that is the question

OpenStudy (dan815):

w o w

OpenStudy (kainui):

:D

OpenStudy (dan815):

learn to medal all nub

OpenStudy (dan815):

:D

OpenStudy (kainui):

pls im not an os hacker

OpenStudy (freckles):

dan likes medal parties

OpenStudy (kainui):

the more you reply here the higher your engagement score gets, probably, right?

OpenStudy (kainui):

I should suspend dan probably right

OpenStudy (freckles):

we can suspend him now and then unsuspend him in 30 cat seconds

OpenStudy (anonymous):

So this has been solved?